# Ideal Triangulations and Once-Punctured Surface Bundles

### v1

@@ 4. The Algorithm
Both algorithms make use of Lemma 3.6. It should be noted in general it may be possible to strategically apply Proposition 3.4 to arrive at an inflation triangulation (not short) where the ideal triangulation obtained by crushing $\partial M$ has the desired property. Currently, the only practical way to check if the fiber is realized as a spun-normal surface is through the use of the computer software tnorm [46] developed by Worden, this software must compute the vertex surfaces for each ideal triangulation. There exist frames for which a desired ideal triangulation is not realized at any intermediate step. One such example appears in Section 5, the three tetrahedra ideal triangulation of the pretzel knot $P(-2,3,7)$ complement in $S^{3}$.
@@ 5. Examples and Questions
The following section contains two examples of the algorithm of Theorem 4.1. A third example is given to illustrate the general absence of fibers among spun-normal surfaces. In each example, Regina [10], SnapPy are utilized via SageMath [40] to construct and manipulate triangulations. Worden’s program ’tnorm’ [46] is used to verify that fibers are represented by embedded spun-normal surfaces.
@@
Using Regina we now enumerate the vertex solutions. Among these surfaces there is a single essential once-punctured torus. This is surface 23 in the enumeration given by Regina version 7.3. By Lemma 2.6 this surface is isotopic to a fiber of the bundle structure on $M_{K}$. The band and branch tetrahedra are shown in Figure 9. After a 2-2 move at the branch-point tetrahedra $\Delta_{3},\Delta_{4}$, we arrive at Case 3 of Theorem 3.11. We then perform a 2-3 move at face (023) of $\Delta_{2}$ and shorten to arrive at the final short inflation in Table 3.
@@
We again enumerate the vertex normal surfaces, finding the fiber, $F$, at surface 43 having no quadrilaterals in the band tetrahedra. We can now create the desired ideal triangulation of $M_{K}$ by crushing along the normal boundary in $\mathcal{T}^{\prime\prime}$. During the crushing processes as implemented by Regina, if $\Delta_{i},\Delta_{j}\in\mathcal{T}^{\prime\prime}$ with $i<j$, then the corresponding ideal tetrahedra in the induced ideal triangulation $\Delta^{*}_{l},\Delta^{*}_{k}\in\mathcal{T}^{**}$ are such that $l<k$. Thus we can take the quadrilateral coordinates of $F$ in $\mathcal{T}^{\prime\prime}$ and remove the corresponding three coordinates for each band and branch tetrahedron. We then have $F$ given as a spun-normal surface in $\mathcal{T}^{**}$ with Q-coordinates:
@@
….: if surface.eulerChar == 0 and surface.hasRealBoundary():
@@
Using Regina, crush Tx along the normal boundary (and finding said normal boundary via Regina). Tnorm then identifies the vertices of Thurston’snormballrepresentedbyembeddedspun-normalsurfacesinRegina’s enumeration at indices 1131 and 219.
@@
TheidealtriangulationcreatedinExample2hashighcomplexitywhencomparedtotheminimaltriangulationofthecomplementofK12n242.Therearelikelylowercomplexityidealtriangulationsinwhichthefiberspun-normalizes.ExaminingtheavailablePachnermovesthatlowercomplexitywefindasingle3-2movethatessentiallyreversesthelastshortening2-3movein“path.”Theeffectintermsofthespun-normalsurfacesistoeliminatethepossibilitythatthefiberspun-normalizes.SearchingthePachnergraphwhilesearchingforatriangulationwithaparticularpropertyisingeneralhard.Forexample,Regina’s strategy for finding a triangulation with fewer tetrahedra is to randomly apply 1-4 moves and check if 3-2 moves appear.
@@
For the complement \$M_K\$ of a hyperbolic knot \$K\$ that has the structure of a fiber bundle over a once-punctured surface \$F\$, does there exist an ideal triangulation admitting an angle structure in which \$F\$ is isotopic to a spun-normal surface?

### v2

@@ 4. The Algorithm
Both algorithms make use of Lemma 3.6. It should be noted in general it may be possible to strategically apply Proposition 3.4 to arrive at an inflation triangulation (not short) where the ideal triangulation obtained by crushing $\partial M$ has the desired property. The current state of the art to verify that the fiber is realized as a spun-normal surface is through the use of the computer software tnorm [45] developed by Worden, this software must compute the vertex surfaces for each ideal triangulation. There exist frames for which a desired ideal triangulation is not realized at any intermediate step. One such example appears in Section 5, the three tetrahedra ideal triangulation of the pretzel knot $P(-2,3,7)$ complement in $S^{3}$.
@@ 5. Examples and Questions
The following section contains two examples of the algorithm of Theorem 4.1. A third example is given to illustrate the general absence of fibers among spun-normal surfaces. In each example, Regina [10] and SnapPy are utilized via SageMath [39] to construct and manipulate triangulations. Worden’s program tnorm [45] is used to verify that fibers are represented by embedded spun-normal surfaces.
@@
Using Regina we now enumerate the vertex solutions. Among these surfaces there is a single essential once-punctured torus. This is surface 23 in the enumeration given by Regina version 7.3. By Lemma 2.6 this surface is isotopic to a fiber of the bundle structure on $M_{K}$. The band and branch tetrahedra are shown in Figure 10. After a 2-2 move at the branch-point tetrahedra $\Delta_{3},\Delta_{4}$, we arrive at Case 3 of Theorem 3.11. We then perform a 2-3 move at face (023) of $\Delta_{2}$ and subsequently shorten to arrive at the final short inflation in Table 3.
@@
We again enumerate the vertex normal surfaces, finding the fiber $F$ at index 43 having no quadrilaterals in the band tetrahedra. We can now create the desired ideal triangulation of $M_{K}$ by crushing along the normal boundary in $\mathcal{T}^{\prime\prime}$. During the crushing processes as implemented by Regina, if $\Delta_{i},\Delta_{j}\in\mathcal{T}^{\prime\prime}$ with $i<j$, then the corresponding ideal tetrahedra in the induced ideal triangulation $\Delta^{*}_{l},\Delta^{*}_{k}\in\mathcal{T}^{**}$ are such that $l<k$. Thus we can take the quadrilateral coordinates of $F$ in $\mathcal{T}^{\prime\prime}$ and remove the corresponding three coordinates for each band and branch tetrahedron. We then have $F$ given as a spun-normal surface in $\mathcal{T}^{**}$ with Q-coordinates:
@@
….: if surface.eulerChar == 0 and not surface.hasRealBoundary():
@@
[…]
@@
For the complement \$M_K\$ of a hyperbolic knot \$K\$ that has the structure of a fiber bundle over a once-punctured surface \$F\$, does there exist an ideal triangulation admitting a strict (taut, semi) angle structure in which \$F\$ is isotopic to a spun-normal surface?